# Ordinal Adaptive Correction: A Data-Centric Approach to Ordinal Image Classification with Noisy Labels

### v1

@@ IV-B Performance under Label Noise
IV-C Analysis of the Correction Mechanism
@@
Crucially, some of these apparent errors (indicated by a small green box) are actually plausible corrections. This happens when the original ”clean” label in the dataset was already incorrect. This result shows that our method can fix not only the noise we add for experiments but also the hidden errors that already exist in the dataset.
@@
TABLE II: MAE and RMSE between the true labels and the labels used for training (Noisy for DLDL-v2, Corrected for ORDAC).
@@
TABLE III: MAE comparison with the CASSOR sample selection method. ”Normal Training” uses all noisy data. ”ORDACC + CASSOR” applies CASSOR’s selection to our corrected dataset.
@@
We also measure the MAE and RMSE between the true labels and the noisy/corrected labels in Table II. While ORDAC sometimes increases the MAE of the labels themselves (e.g., for DR at $\tau=0.2$), it consistently improves the final model’s performance on the test set (Table I). This indicates that simply measuring label correctness is insufficient; the ultimate goal is to produce a more generalizable model, which ORDAC achieves. The improved model performance suggests that even when a label is not corrected perfectly, it is often moved closer to the true value, or its associated uncertainty is adjusted appropriately, leading to better overall training dynamics.
IV-D Comparison with State-of-the-Art Sample Selection
We compare our correction-based approach with CASSOR [7], a state-of-the-art sample selection method for noisy ordinal regression. To ensure a fair comparison, we use their official implementation and train it with the same ResNet-50 backbone. We evaluate training with CASSOR’s selected clean subset versus training with our corrected labels (ORDACC and ORDACR).
The results in Table III show that our correction-based methods generally outperform sample selection, especially as the noise rate increases. On the DR dataset with 40% noise, ORDACR achieves an MAE of 0.7435, significantly better than CASSOR’s 0.8240. This suggests that correcting noisy labels is a more data-efficient strategy than simply discarding them. An interesting case is Adience at $\tau=0.4$, where CASSOR performs best. This may indicate that at very high noise levels, aggressive filtering can be more effective than attempting to correct highly unreliable labels. We also test a hybrid approach (ORDACC + CASSOR), which yields strong results, suggesting that correction and selection are complementary.
IV-E Ablation Studies
IV-E1 Effect of Class-wise Prediction Debiasing
We evaluate the importance of the class-wise debiasing step described in Section III-B. Table IV shows that removing this step significantly degrades performance, especially at higher noise rates. As shown in Figure 5, without debiasing, the correction process develops a strong bias towards the majority middle class, drastically depleting the minority classes at the ends of the spectrum. The debiasing step successfully mitigates this, ensuring a more balanced and accurate correction process.
TABLE IV: Effect of removing the class-wise prediction debiasing step on the Adience dataset.
@@ IV-E1 Effect of Class-wise Prediction Debiasing
IV-E2 Hyperparameter Analysis
We analyzed the sensitivity to the base correction rates, $\alpha_{\text{base}}$ and $\beta_{\text{base}}$, on Adience with $\tau=0.4$. The results in Table V show that the best performance is achieved with a small $\alpha_{\text{base}}$ (0.2) and a large $\beta_{\text{base}}$ (0.8). This aligns with our intuition: the model should be aggressive in correcting the label’s mean value ($\mu$) but conservative and stable when updating its uncertainty ($\sigma$). We also found that an initial standard deviation ($std_{\text{init}}$) of 0.75 yielded the best results, providing a good balance for initial uncertainty representation.
TABLE V: Hyperparameter tuning results (MAE) for $\alpha_{\text{base}}$ and $\beta_{\text{base}}$ on Adience ($\tau=0.4$).

### v2

@@ 4.2 Performance under Label Noise
4.3 Analysis of the Correction Mechanism
@@
Crucially, some of these apparent errors (indicated by a small green box) are actually plausible corrections. This happens when the original "clean" label in the dataset was already incorrect. This result shows that our method can fix not only the noise we add for experiments but also the hidden errors that already exist in the dataset.
@@
Table 2: MAE and RMSE between the true labels and the labels used for training (Noisy for DLDL-v2, Corrected for ORDAC).
@@
Table 3: MAE comparison with the CASSOR sample selection method. "Normal Training" uses all noisy data. "ORDACC + CASSOR" applies CASSOR’s selection to our corrected dataset.
@@
We also measure the MAE and RMSE between the true labels and the noisy/corrected labels in Table 2. While ORDAC sometimes increases the MAE of the labels themselves (e.g., for DR at $\tau=0.2$), it consistently improves the final model’s performance on the test set (Table 1). This indicates that simply measuring label correctness is insufficient; the ultimate goal is to produce a more generalizable model, which ORDAC achieves. The improved model performance suggests that even when a label is not corrected perfectly, it is often moved closer to the true value, or its associated uncertainty is adjusted appropriately, leading to better overall training dynamics.
4.4 Comparison with State-of-the-Art Sample Selection
We compare our correction-based approach with CASSOR [yuan2023cassor], a state-of-the-art sample selection method for noisy ordinal regression. To ensure a fair comparison, we use their official implementation and train it with the same ResNet-50 backbone. We evaluate training with CASSOR’s selected clean subset versus training with our corrected labels (ORDACC and ORDACR).
The results in Table 3 show that our correction-based methods generally outperform sample selection, especially as the noise rate increases. On the DR dataset with 40% noise, ORDACR achieves an MAE of 0.7435, significantly better than CASSOR’s 0.8240. This suggests that correcting noisy labels is a more data-efficient strategy than simply discarding them. An interesting case is Adience at $\tau=0.4$, where CASSOR performs best. This may indicate that at very high noise levels, aggressive filtering can be more effective than attempting to correct highly unreliable labels. We also test a hybrid approach (ORDACC + CASSOR), which yields strong results, suggesting that correction and selection are complementary.
4.5 Ablation Studies
4.5.1 Effect of Class-wise Prediction Debiasing
We evaluate the importance of the class-wise debiasing step described in Section 3.2. Table 4 shows that removing this step significantly degrades performance, especially at higher noise rates. As shown in Figure 5, without debiasing, the correction process develops a strong bias towards the majority middle class, drastically depleting the minority classes at the ends of the spectrum. The debiasing step successfully mitigates this, ensuring a more balanced and accurate correction process.
Table 4: Effect of removing the class-wise prediction debiasing step on the Adience dataset.
@@ 4.5.1 Effect of Class-wise Prediction Debiasing
4.5.2 Hyperparameter Analysis
We analyzed the sensitivity to the base correction rates, $\alpha_{\text{base}}$ and $\beta_{\text{base}}$, on Adience with $\tau=0.4$. The results in Table 5 show that the best performance is achieved with a small $\alpha_{\text{base}}$ (0.2) and a large $\beta_{\text{base}}$ (0.8). This aligns with our intuition: the model should be aggressive in correcting the label’s mean value ($\mu$) but conservative and stable when updating its uncertainty ($\sigma$). We also found that an initial standard deviation ($std_{\text{init}}$) of 0.75 yielded the best results, providing a good balance for initial uncertainty representation.
Table 5: Hyperparameter tuning results (MAE) for $\alpha_{\text{base}}$ and $\beta_{\text{base}}$ on Adience ($\tau=0.4$).